From: Frank Geurts (geurts_at_rice.edu)
Date: Mon Mar 17 2003 - 11:32:51 EST
called SCR: they are experiencing some temperature problems right now
(which means they'll start shutting down electronics in the WAH any time
soon) and apparently cannot do any pedestal test runs. I'll contact them
later today for our test.
-frank
Frank Geurts wrote:
> yep, tof ran only for a brief period between 23h11 and 2h18 this morning
> .. like before no apparent problems from the daq-operation.
>
> More later.
>
> -frank
>
>
>
>
> Tonko A. Ljubicic wrote:
>
>> On Mon, 17 Mar 2003, W.J. Llope wrote:
>>
>>
>>> ...ran fine. data looks A.O.K.....
>>>
>>> looks like eleanor's reseating of a lemo cable at the TOF TCD
>>> did the trick... thanks again eleanor..
>>
>>
>>
>> Hm, according to the shift-leader, it did NOT run fine and as such is
>> NOT in the run... I don't know anything more...
>>
>> -- Tonko
>>
>>
>> --
>> Tonko (Ante) Ljubicic Email: tonko_at_bnl.gov
>> Bldg. 510A Tel: +1-631-344-7346
>> Brookhaven National Lab., P.O. Box 5000 Fax: +1-631-344-4206
>> Upton, NY, 11973
>>
>
This archive was generated by hypermail 2.1.4 : Thu Jul 24 2003 - 00:39:38 EDT